I need to find all of the positive values for k for which each of the following can be factored.
xsquared+x-k

Answer by stanbon(48564)

find all of the positive values for k for which each of the following can be factored.
xsquared+x-k
-----------
x^2+x-k
a=1,b=1,c=-k
-----------------
Comments: The quadratic will be factorable with Real Number factors
if b^2-4ac is greater than or equal to zero.
---------------
Consider b^2-4ac >=0
1^2-4(1)(-k)>=0
1+4k>=0
4k>=-1
k>=-1/4
Factorable with Real Number factors when k>=(-1/4)
----------
If you want k to be positive, let k>0.
===========
Cheers,
Stan H.
-------------------